OpenStudy (nerdy_3000):

4sqrt 10x^5 y^8 z^10

OpenStudy (solomonzelman):

\[\Large 4\sqrt {10x^5 y^8 z^{10}}\]

OpenStudy (solomonzelman):

This ?

OpenStudy (nerdy_3000):

Yes how can I simplify this?

OpenStudy (solomonzelman):

Tips:\[x^{5}=x^4 \times x=(x^2)^2 \times x\]\[y^8=(y^4)^2 \]\[z^{10}=(z^{5})^2\]

OpenStudy (nerdy_3000):

Thank you!

OpenStudy (solomonzelman):

Anytime !